The width of a rectangle is 5 meters less than its length. The area is 84 square meters. Find the length of the rectangle.
Answer by AHBK(13)

let width be W , and length be L.
we have W= L-5
and we have area= 84 meter square .
so W multiplied by length L is equal to 84
(W)(L)=84
substitute L-5 instead of W
which gives us that (L-5)(L)=84 so,
L^2 -5L =84 then
L^2 -5L -84 =0 which gives us a second degree equation
so by solving it we get
L=12m or L= -7m
and since we r talking about length then it should be positive .
then L= -7m is rejected and L= 12m is accepted .
so length of rectangle is 12m
note : ()() means they r multiplied by each other